OpenStudy (anonymous):

Which of the following describes the transformation of the graph y = x^2 in graphing y = -x^2 - 5?

OpenStudy (anonymous):

@angelina22309 will u help me

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (callisto):

Start from y = x^2 y = -x^2 means to reflect y=x^2 about the x axis. y = -x^2 -5 means to reflect y=x^2 about the x-axis, then move the graph downwards by 5 units. Actually, what is your question?

OpenStudy (anonymous):

u got it and that's I understand it

OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (0, 3)?

OpenStudy (callisto):

1. Please start a new post for a new question. 2. Please specify which part you don't understand. 3. That is not a complete question.

OpenStudy (anonymous):

okay and ya it was a complete question and I do understand it